how do i solve 2x-3y=6
2x-y=7
substitution method linear systems

Answer by jim_thompson5910(35256)

Hint:

Solve the second equation for y

2x-y=7

-y = -2x+7

y = 2x-7

Then plug this into the first equation

2x-3y=6

2x-3(2x-7) = 6

2x-6x+21 = 6

-4x+21 = 6

I'll let you finish. Solve for x. Then use the solution for x to find the solution for y.